Calculate range and its coefficient from the following data:


Price of sold per 1 gm from Monday to Saturday


Monday Tuesday Wednesday Thursday Friday Saturday


902 910 914 908 922 920

1
Expert's answer
2021-12-29T16:32:22-0500

Range: "922-902=20."

Coefficient of range: "\frac{922-902}{922+902}=\frac{20}{1824}\approx 0.0110."